### v1

# MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning

Abstract
Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computation and memory cost. Previous LoRA-based approaches initialize the low-rank matrices with gaussian distribution and zero values, while keeping the original weight matrices frozen. However, the trainable model parameters optimized in an unguided subspace might have interference with the well-learned subspace of the pretrained weight matrix. In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principle singular components frozen. It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principle matrix contains important knowledge. The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principle matrix, thus the pretrained knowledge is expected to be well preserved. During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the finetuning dataset. Extensive experiments on commonsense reasoning, math reasoning and instruction following benchmarks present the superior performance of our method.
1 Introduction
Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024) and question answering (Ivison et al., 2023). These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks. Fully finetuning is commonly employed to unlock the complete potential of LLMs, however, optimizing all model parameters necessitates substantial and restrictive computational resources (Touvron et al., 2023; Jiang et al., 2023), which hampers the utilization of LLMs across diverse scenarios.
Parameter-efficient finetuning (Hu et al., 2021, 2023b; Liu et al., 2022; Zhang et al., 2024; Liu et al., 2024a, PEFT) aims at reducing the computational and GPU memory cost for finetuning of pretrained models. Low-rank adaptation (Hu et al., 2021, LoRA) is one of the most widely used PEFT methods for LLM finetuning. It assumes the change of linear model weights to be low-rank (Li et al., 2018; Aghajanyan et al., 2021). For each selected weight matrix, it only updates two low-rank matrices while keeping the pretrained weight frozen. During inference, the low-rank matrices are merged into pretrained linear weights, thus no additional computation or memory cost is introduced. Recently, researchers explore different LoRA-based variants for efficient LLM finetuning (Zhang et al., 2023; Pan et al., 2024; Kopiczko et al., 2024; Liu et al., 2024a; Meng et al., 2024). However, most existing LoRA-based works randomly initialize the low-rank matrices and optimize the trainable parameters in an unguided subspace. We suspect this strategy may override important pretrained features, thus influencing the performance of low-rank adaptation methods (Dou et al., 2024).
In this paper, we propose Minor singular component based Low Rank Adaptation (MiLoRA) for efficient LLM finetuning. MiLoRA has a similar framework as vanilla LoRA, but employs a different initialization schedule. Specifically, a weight matrix ${\bm{W}}$ is decomposed with singular value decomposition (SVD) algorithm. Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principle matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values. We argue that the principle matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information. It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that using principle low-rank approximation matrices can achieve comparable or even better performance to full parameters.
Motivated by these observations, we keep the principle matrix ${\bm{W}}_{p}$ frozen and adapt the minor singular components during finetuning. The low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ in LoRA framework are initialized with the minor matrix ${\bm{W}}_{m}$. Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principle matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge. To maintain the capability of the pretrained model at the start of finetuning, vanilla LoRA explicitly initializes ${\bm{B}}$ with zeros. In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals to the frozen principle part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$. We conduct extensive experiments on commonsense reasoning, math reasoning and instruction following benchmarks. The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+1.92 on LLaMA2-7B), and instruction following (+1.4 on LLaMA2-7B). [^1] Our code and model will be made public.
2 Preliminaries
Singualr Value Decomposition
Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition (SVD) is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$. The columns of ${\bm{U}}$ are the left singular vectors, and the columns of ${\bm{V}}$ are the right singular vectors. The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order. Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}=\sum_{i=1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$ (1)
where $u_{i}$ and $v_{i}$ are the $i^{\mathrm{t}h}$ column of ${\bm{U}}$ and ${\bm{V}}$, respectively.
Low-Rank Adaptation
The low-rank adaptation method (Hu et al., 2021, LoRA) assumes the updates of linear weight ${\bm{W}}\in\mathbb{R}^{m\times n}$ to be low-rank, thus models the changes with two trainable low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times n}$ and ${\bm{B}}\in\mathbb{R}^{m\times r}$. The weight matrix can be decomposed as
${\bm{W}}={\bm{W}}^{(0)}+\Delta{\bm{W}}={\bm{W}}^{(0)}+\frac{\alpha}{r}{\bm{B}}{\bm{A}},$ (2)
where ${\bm{W}}^{(0)}$ and $\Delta{\bm{W}}$ refer to the pretrained weight and weight change, respectively. The $\alpha$ and $r$ are hyperparameters of scaling-factor and LoRA rank ($r\ll\min(m,n)$). During finetuning, the pretrained matrix ${\bm{W}}^{(0)}$ is kept frozen. It significantly diminishes the number of trainable parameters as both ${\bm{A}}$ and ${\bm{B}}$ matrices are low-rank. The ${\bm{B}}$ matrix is initialized with zero while ${\bm{A}}$ matrix adopts a random gaussian distribution with zero mean value. This initialization strategy ensures the $\Delta{\bm{W}}=0$ at the beginning of training. The LoRA method only modifies the linear matrices in the Transformer model. The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with fully finetuning. However, vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized. This might potentially detract from the pretrained knowledge encoded in the principal weight matrix.
Figure 1: (a) MiLoRA method splits the pretrained linear weight matrix into two parts, the minor singular component (${\bm{U}}_{m},{\bm{V}}_{m},{\bm{\Sigma}}_{m}$) is used to initialize the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ in the LoRA framework. (b) During finetuning, only the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ are updated while the principle matrix ${\bm{W}}_{p}$ is frozen, which is similar as the practice of LoRA method.
3 Methodology
The proposed MiLoRA is a simple yet effective PEFT approach that selects the minor singular components of the weight matrices for optimization. As observed in LASER (Sharma et al., 2024), the minor singular components of weight matrices contain noisy or long-tail information, while the principle singular components are responsible for important features across tasks. Therefore, MiLoRA is designed to effectively learn from the finetuning dataset by adapting minor singular components, while maintaining the knowledge encoded in the pretrained model. On the contrary, vanilla LoRA fails to constrain the optimization subspace as its low-rank matrices are randomly initialized. Specifically, suppose $m\leq n$, MiLoRA splits each linear weight matrix into two matrices according to the corresponding singular values: principal matrix ${\bm{W}}_{p}$ and minor matrix ${\bm{W}}_{m}$ (see Figure 1):
${\bm{W}}={\bm{W}}_{p}+{\bm{W}}_{m}=\sum_{i=1}^{m-r}\sigma_{i}u_{i}v_{i}^{\top}+\sum_{i=m-r+1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$ (3)
where the singular values are in descending order ($\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}\geq 0$), and the hyperparameter $r$ is the number of minor singular values in the ${\bm{W}}_{m}$ matrix.
The decomposition can be also reformulated in a matrix form. The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principle and minor singular values, respectively. The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated in a similar way. The SVD of ${\bm{W}}$ can be expressed as:
${\bm{W}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}={\bm{U}}_{p}{\bm{\Sigma}}_{p}{\bm{V}}_{p}^{\top}+{\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}={\bm{W}}_{p}+{\bm{W}}_{m}$ (4)
During the finetuning process, instead of freezing the entire pretrained weight matrix, we just keep the principal singular components ${\bm{W}}_{p}$ fixed to preserve the pretrained knowledge. The minor matrix ${\bm{W}}_{m}$ is used to initialize low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ (see Figure 1):
${\bm{W}}_{m}={\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}=({\bm{U}}_{m}\sqrt{{\bm{\Sigma}}_{m}})(\sqrt{{\bm{\Sigma}}_{m}}{\bm{V}}_{m}^{\top})={\bm{B}}_{m}{\bm{A}}_{m}$ (5)
This strategy has two benefits: 1) It encourages the model to learn in the less-optimized subspace spanned by the minor singular vectors, thus reducing the interference with the well-learned pretrained knowledge encoded in the principle singualr components. 2) Unlike vanilla LoRA which requires tuning the scaling factor $\alpha$ and the initialization hyperparameters, ours does not require any hyperparameter except the rank $r$. To make our method simple and training-efficient, we do not incorporate the regularization loss which imposes strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$. Our experiment results show that our scheme works quite well in balancing learning from finetuning dataset and preverving pretrained knowledge.
4 Experiments
To evaluate our method against other PEFTs, we conduct experiments across three diverse tasks covering more than 10 datasets. Our objective is to illustrate the effectiveness of our method across various scenarios. We benchmark against LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024).
  • LoRA (Hu et al., 2021) reparameterizes the weight update $\Delta{\bm{W}}$ with two trainable low-rank matrices ${\bm{A}}$ and ${\bm{B}}$, while freezing the pretrained weight ${\bm{W}}$. They use a random Gaussian initialization for A and zero for B, so $\Delta{\bm{W}}={\bm{B}}{\bm{A}}$ is zero at the beginning of training.
  • PiSSA (Meng et al., 2024) shares the same framework as LoRA, but it employs a more sophisticated initialization approach. Specifically, they initialize ${\bm{A}}$ and ${\bm{B}}$ with principal singular values and singular vectors of the pre-trained weight ${\bm{W}}$. Given that the principal singular values and vectors capture the essence of a matrix, PiSSA is hoped to better approximate the outcomes of full-parameter finetuning by changing the essential parts while freezing the “noisy” parts.
The MiLoRA method optimizes the low-rank matrices within a less-optimized subspace, which can be combined with other LoRA-based PEFT methods (Liu et al., 2024a; Zhang et al., 2023; Pan et al., 2024). We leave the exploration of combination MiLoRA with LoRA variants as future work.
4.1 Commonsense reasoning
Setup
We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023a). Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The task is formulated as a multiple-choice problem. The best checkpoint is chosen by their validation set loss. We report accuracy (%) for all datasets. We use the implementation of LLM-Adapters (Hu et al., 2023b) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters.
Hyperparameter
We use the same hyperparameter configurations as Hu et al. (2023a) without tuning for all methods. Details can be found in Table 8 of Appendix A.2. We denote this hyperparameter setup as LLM-Adapters.
Results
Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks. Performance results of baseline methods are taken from Liu et al. (2024a). Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA2-7B LoRA 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA 67.6 83.8 80.1 88.2 82.0 82.8 68.8 80.6 79.2
LLaMA3-8B LoRA 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PiSSA 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
We report results in Table 1. We also include the ChatGPT baseline reported in Liu et al. (2024a) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022). MiLoRA consistently surpasses all baseline methods across both LLaMA2-7B and LLaMA3-8B models. On LLaMA2-7B, MiLoRA outperforms LoRA and PiSSA by 1.6 and 5.4 average accuracy scores. On LLaMA3-8B, MiLoRA exceeds LoRA and PiSSA by 1.1 and 6.5 average accuracy scores. These outcomes suggest the effectiveness of MiLoRA as a PEFT method.
4.2 Math reasoning
Setup
We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH(Hendrycks et al., 2021). We use test set of GSM8K and MATH for evaluation and report results on the last checkpoint. For each test set, we report the exact match ratio against the ground truth. We use the implementation of PiSSA(Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA.
Hyperparameter
As MetaMathQA has a large training dataset, we set the rank to 64. For all other hyperparameters, we follow the llm-adapter hyperparameter setup shown in Table 8 of Appendix A.2. We limit all samples to a maximum length of 2048 tokens.
Results
Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH. Results are cited from Yu et al. (2023).
Method GSM8K MATH Avg.
Full FT 66.5 19.8 43.2
LoRA 60.58 16.88 38.73
PiSSA 58.23 15.84 37.04
MiLoRA 63.53 17.76 40.65
We report results in Table 2. We also include the full finetuning results from (Yu et al., 2023) as a reference. We can see that MiLoRA achieves the best performance on LLaMA2-7B. However, it still underperforms full finetuning by an average of 2.55 EM score. This suggests that PEFT methods still have room for improvement to fully match the performance of full finetuning.
4.3 Instruction-following
Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Performance results of baseline methods are taken from Ivison et al. (2023).
Method Win-rate
GPT-4-0613 91.2
GPT-3.5-turbo-0613 91.8
LLaMA2 Chat 7B 87.3
LLaMA2 7B + LoRA 85.9
LLaMA2 7B + PiSSA 85.6
LLaMA2 7B + MiLoRA 87.3
Setup
We follow the experiment setup in Wu et al. (2024a) and Wu et al. (2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023). Following previous works, we use Alpaca-Eval v1.0  (Li et al., 2023) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024). We implement our method using open-instruct [^4] https://github.com/allenai/open-instruct.
Hyperparameter
We use the LLM-Adapters hyperparameter setup for all PEFT methods. We truncate samples to a maximum of 768 tokens. During evaluation, we use the last checkpoint with a greedy decoding strategy and a maximum token number of 2048. To avoid the issue of repetition, we set the repetition penalty to 1.1 and set the no repeat n-gram size to 5, following Wu et al. (2024a) and Wu et al. (2024b).
Results
Table 3 presents the win-rate of various PEFT methods. We include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2 Chat 7B as a reference, which have win-rate of 91.2%, 91.8%, and 87.3% respectively. Our method, LLaMA2-7B + MiLoRA outperforms the other two PEFT methods and matches LLaMA2 Chat 7B’s performance, indicating that MiLoRA is a highly effective finetuning method.
5 Understanding MiLoRA
In this section, we conduct experiments to further understand MiLoRA. For all analysis experiments that do not involve finetuning, we use the LLaMA2-7B model finetuned on MetaMathQA in our main experiments (see Section 4.2). For any experiments that require finetuning except for Section 5.5, we finetune the LLaMA2-7B model on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512. This configuration is chosen to minimize computational costs.
5.1 How do different rank settings influence MiLoRA?
Figure 2: Average score of finetuned LLaMA2-7B model on math reasoning across different ranks.
In Figure 2, we explore the impact of various rank configurations by adjusting $r$ within the set $\{16,32,64,128\}$ and assessing the average accuracy of the finetuned LLaMA2-7B model on math reasoning tasks. Detailed numbers are presented in Table 10 of Appendix B.1. As can be seen, MiLoRA with rank 64 achieves the best performance across all methods and ranks. MiLoRA consistently surpasses PiSSA across all rank configures. However, when compared with LoRA, MiLoRA performs better only at large ranks. This is likely due to the fact that MiLoRA makes fewer changes to the pre-trained parameters compared to LoRA (see detailed discussion in Section 5.3 and 5.4). As a result, when the rank is relatively low, MiLoRA struggles to fully digest the finetuning dataset.
5.2 Is minor singular component adaptation important?
To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model. Specifically, we initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and evaluate the results on the GSM8K and MATH datasets.
Table 4: Performance of MiLoRA when initializing with principal, random sampled, and minor singular components.
Method Principle Random Minor
GSM8K 60.65 63.23 63.99
MATH 14.60 15.54 16.14
The results, shown in Table 4, demonstrate that initializing with the minor singular components achieves the best performance across both evaluation datasets. This underscores the importance of adapting the minor singular components during the finetuning process. Concurrent with our work, Meng et al. (2024) propose PiSSA, a similar approach to MiLoRA but adapts the principle singular components. We will provide a detailed comparison between MiLoRA and PiSSA in Section 5.5.
5.3 How does the matrix update $\Delta W$ compare to $W$?
We investigate the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ of different methods, where $\Delta{\bm{W}}$ is the difference between the finetuned and pre-trained weights. Following the analysis method in Hu et al. (2021), we project ${\bm{W}}$ and $\Delta{\bm{W}}$ onto the $r$-dimensional subspace of $\Delta{\bm{W}}$ by computing ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ and ${\bm{U}}^{T}\Delta{\bm{W}}{\bm{V}}$, with ${\bm{U}}$/${\bm{V}}$ being the left/right top $r$ singular-vector matrix of $\Delta{\bm{W}}$. As a comparison, we also compute ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ by replacing ${\bm{U}}$, ${\bm{V}}$ with the top $r$ singular vectors of ${\bm{W}}$ or a random matrix. We use the Frobenius norm as the measurement of magnitude. The results are shown in Table 5. Following Hu et al. (2021), we use the query weight in the middle layer (16th) of the model for analysis. We make two key observations. First, $\Delta{\bm{W}}$ of LoRA and MiLoRA has a similar correlation with ${\bm{W}}$, both slightly stronger compared to a random matrix, indicating that they amplify directions that are not emphasized in ${\bm{W}}$. In contrast, $\Delta{\bm{W}}$ of PiSSA shows a strong correlation with ${\bm{W}}$. This is attributed to PiSSA’s direct optimization of the principal singular components. Second, the amplification factor varies significantly across different methods. Compared to LoRA, we find the amplification factor of MiLoRA is significantly smaller. This indicates that MiLoRA has a reduced impact on features already present in ${\bm{W}}$. See Table 11 of Appendix B.2 for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
Table 5: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs. The weight matrices are taken from the query projection of the 16th layer of LLaMA2-7B.
$\left\|{\bm{W}}\right\|_{F}=97.36$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
$\left\|{\bm{U}}^{\top}{\bm{W}}{\bm{V}}\right\|_{F}=$ 43.64 1.50 1.82 17.57 1.86
$\left\|{\bm{U}}^{\top}\Delta{\bm{W}}{\bm{V}}\right\|_{F}=$ 68.18 55.79 44.95
Amplification factor - - 37.46 3.18 24.17
To further analyze the characteristics of $\Delta{\bm{W}}$ for different methods, we measure the normalized subspace similarity (Hu et al., 2021) between the subspaces spanned by top $r$ singular vectors of $\Delta{\bm{W}}$ and the subspace spanned by top $r$ singular vectors, bottom $r$ singular vectors, and a random $r$ singular vectors of ${\bm{W}}$. We define the normalized similarity metric based on the Grassmann distance following Hu et al. (2021) (See Appendix A.1 for more details). The results are presented in Figure 3. From the left figure, it is evident that PiSSA focuses on optimizing features associated with the top singular vectors, whereas LoRA and MiLoRA do not significantly optimize these features. MiLoRA primarily optimizes features related to the bottom singular vectors. The right figures further show that MiLoRA places greater emphasis on features associated with the bottom singular vectors of the MLP layers.
Figure 3: Left: The subspace similarity averaged by all modules in a layer. Right: The subspace similarity averaged by all layers for a specific module type.
5.4 Does MiLoRA forget less than baselines?
Table 6: Forgetting loss of various PEFT methods.
Method LoRA PiSSA MiLoRA
Forgetting loss 3.24 6.07 2.54
A hypothesis explanation for why MiLoRA outperforms baselines is that pre-trained knowledge is more retained. To test this hypothesis, we follow Kalajdzievski (2024) and use cross-entropy as the metric for measuring forgetting. This is the usual next token prediction loss used when training LLMs, except that the target next token is replaced by the distribution predicted by the pre-trained base model. We evaluate the forgetting metric on LLaMA2-7B model finetuned in Section 4.2 using the WikiText-103 test dataset following (Kalajdzievski, 2024). As shown in Table 6, MiLoRA exhibits the lowest forgetting loss, which is consistent with our hypothesis that MiLoRA makes the least modification to pre-trained knowledge. In contrast, the loss of PiSSA is significantly higher than that of the other methods.
5.5 Comparison between MiLoRA and PiSSA
Concurrent with our research, Meng et al. (2024) have recently proposed a low-rank adaptation method called PiSSA. At first glance, PiSSA may appear to be quite similar to MiLoRA but adapts the principle singular components. However, there are fundamental differences between PiSSA and MiLoRA that are worth highlighting:
  • Motivation: PiSSA is designed to approximate full finetuning by adapting the principal singular components, which are believed to capture the essence of the weight matrices. In contrast, our method MiLoRA aims to adapt to new tasks while maximally retaining the base model’s knowledge. To achieve this, we instead finetune the minor singular components of the weight matrices, which are less important for the weight matrices.
  • Performance: The PiSSA paper claims that PiSSA outperforms both LoRA and adapting minor singular components. However, our investigation suggests that this claim is likely due to the specific hyperparameters used in their experiments. To verify this, we replicate their experimental setup using the same MetaMathQA 100K dataset and compare the performance of PiSSA, LoRA, and MiLoRA under both PiSSA’s hyperparameters and our own. Specifically, we choose a rank of 128 where PiSSA obtains the best performance in their paper and finetunes for 1 epoch for all setups following PiSSA paper. The results are summarized in Table 7. We have several observations. First, under the PiSSA hyperparameter setup, PiSSA does indeed outperform both LoRA and MiLoRA, consistent with the claims made in the original PiSSA paper. Second, we find that all methods perform better in our hyperparameter setup, suggesting that our hyperparameter configuration is more effective. Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.39 and 4.27 average scores, respectively. In summary, when using an appropriate hyperparameter configuration, our proposed MiLoRA method achieves markedly superior performance compared to PiSSA.
We further investigate why the hyperparameter configuration in PiSSA paper is suboptimal. This configuration, as shown Table 9 of Appendix A.3, is from the full finetuning hyperparameters of Alpaca (Taori et al., 2023). As can be seen, it has a combination of small learning rate (2e-5) and large batch size (128). This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023). This problem becomes more severe as PiSSA paper only trains for a single epoch for all experiments.
Figure 4: Validation loss curve of different PEFTs under PiSSA hyperparameter configurations. LoRA and MiLoRA are under-fitting when training with 1 epoch (781 steps).
To validate our assumption, we extract another 120 samples from MetaMathQA to serve as a validation set and then train LoRA, PiSSA, and MiLoRA under the same hyperparameter configuration except for 10 epochs each and recorded their respective validation loss curves. From the results shown in Figure 4, we make two key observations. First, LoRA and MiLoRA are much under-fitting when training with only 1 epoch (781 steps). However, with sufficient training, their validation loss can converge to the same level as that of PiSSA. Second, PiSSA converges faster than the other two methods but quickly overfits. This might be attributed to that PiSSA directly finetunes the essence of weight via the top singular vectors and values. We evaluate the best checkpoint of each methods selected by the validation set, and find that LoRA performs best, while the other two perform similarly (see Table 12 in Appendix B.3), which again confirms that LoRA and MiLoRA are observed worse than PiSSA in the PiSSA paper because they are underfitting.
Table 7: Math reasoning evaluation results for LLaMA2-7B with PiSSA hyperparameters and our (LLM-Adapters) hyperparameters.
Hyp. Setup PEFT GSM8K MATH Avg.
PiSSA LoRA 41.47 5.76 23.6
PiSSA 51.25 7.58 29.4
MiLoRA 40.03 5.18 22.6
LLM-Adapters LoRA 56.56 10.84 33.70
PiSSA 51.25 10.38 30.82
MiLoRA 58.61 11.56 35.09
6 Related Work
Parameter-Efficient Finetuning of LLMs.
After pretraining, large language models are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023). Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of fully finetuning. Previous PEFT methods can be grouped into three lines. The first group is adapter-based, which inserts additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b). However, adapter-based methods introduce additional computation costs during inference. The second category is LoRA-based (Hu et al., 2021; Liu et al., 2024a; Zhang et al., 2023), which models the changes of selected linear weights as low-rank matrices. During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen. The third group is prompt-based (Lester et al., 2021; Li & Liang, 2021; Liu et al., 2022) which adds additional soft prompts to the input tokens. During training, only soft prompts are updated while the pretrained model parameters are fixed. Among these PEFT methods, the LoRA-based approaches are most widely used for LLM finetuning as it does not introduce computation overhead during inference while approximating the performance of fully finetuning.
LoRA and Its Variants
LoRA (Hu et al., 2021) reparameterizes the weight update with two trainable low-rank matrices, while freezing the pretrained weights. With this lightweight decomposition, LoRA reduces storage and task-switching overhead by sharing the pretrained models across multiple tasks. Theoretically, the linear design of LoRA enables seamless merging of trainable matrices with frozen weights, thereby avoiding extra inference latency. Since then, researchers explore and propose different LoRA variants for PEFT. AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024b), adaptively determine the rank of LoRA module in each weight matrix according to the importance score. The pruned ranks are compensated on the more important positions. The rsLoRA (Kalajdzievski, 2023) modifies the LoRA with the appropriate scaling factor to improve the performance of large ranks. The DoRA (Liu et al., 2024a) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance. VeRA (Kopiczko et al., 2024) reduces the number of trainable parameters compared to LoRA by sharing a pair of low-rank matrices across all layers. Concurrent with our work, PISSA (Meng et al., 2024) proposes to only update the principle singular components. However, their conclusion is empirically problematic due to the sub-optimal hyperparameters.
7 Conclusion
In this paper, we introduce MiLoRA, a simple yet effective low-rank adaption method for LLMs. MiLoRA effectively learns on finetuning tasks while better preserving the pretrained knowledge by adapting the minor singular components of pretrained weight matrices. We investigate the effectiveness of MiLoRA on a wide range of LLM evaluation benchmarks, including commonsense reasoning, math reasoning and instruction-following. Experiment results demonstrate that MiLoRA consistently outperforms LoRA and PiSSA without sacrificing training or inference efficiency. We hope that our work will inspire future research on parameter-efficient finetuning of LLMs.
Appendix A Detailed Experiment Setups
A.1 Similarity metric between subspace
Following Hu et al. (2021), we use the measure $\phi({\bm{A}},{\bm{B}})=\frac{\left\|{\bm{A}}^{\top}{\bm{B}}\right\|_{F}^{2}}{r}$ to measure the similarity between two column orthonormal matrices ${\bm{A}},{\bm{B}}\in\mathbb{R}^{d\times r}$, where $F$ donotes the Frobenius norm. The value of $\phi({\bm{A}},{\bm{B}})$ ranges from 0 to 1, where 1 indicates complete overlap and 0 indicates complete separation.
Given two matrices $\Delta{\bm{W}}$ and ${\bm{W}}$, we extract $r$ left singular vectors from each to form the subspace matrices for them, which are denoted as $\Delta{\bm{W}}_{s}$ and ${\bm{W}}_{s}$ , and then the subspace similarity is computed using $\phi(\Delta{\bm{W}}_{s},{\bm{W}}_{s})$.
A.2 Our hyperparameter setup
Table 8 shows our detailed hyperparameters. This setup follows LLM-adapters (Hu et al., 2023b), therefore we denote it as LLM-adapters setup.
Table 8: Our hyperparameter configuration on the commonsense reasoning (ComR), math reasoning (MathR) and instruction-following (InsF) tasks.
Hyperparameters ComR MathR InsF
Rank r 32 64 32
$\alpha$ of LoRA 64 128 64
$\alpha$ of PiSSA/MiLoRA 32 64 32
Dropout 0.05
Optimizer AdamW
LR 3e-4
LR Scheduler Linear
Batch size 16
Warmup Steps 100
Epochs 3
Placement query, key, value, MLP up, MLP down
A.3 PiSSA hyperparameter setup
Table 9 shows the detailed hyperparameters from the PiSSA paper, which we denoted as PiSSA hyperparameter setup.
Table 9: Hyperparameter configuration from the PiSSA paper.
PiSSA hyperparameters
$\alpha$ Same as rank r
Dropout 0.0
Optimizer AdamW
LR 2e-5
LR Scheduler cosine
Batch Size 128
Warmup Ratio 0.03
Epochs 1
Placement query, key, value, output, gate, MLP up, MLP down
A.4 Computation resource
All of our experiments are conducted on 8 NVIDIA L40 GPUs.
Appendix B More Experiment Results
B.1 Detailed results on math reasoning across ranks
Detailed EM scores of MiLoRA and baselines on math reasoning by adjusting rank $r$ within the set $\{16,32,64,128\}$.
Table 10: Math reasoning evaluation results for LLaMA2-7B across different ranks.
Method Dataset 16 32 64 128
LoRA GSM8K 62.93 63.38 63.61 59.59
MATH 15.52 15.56 15.28 15.28
PiSSA GSM8K 61.56 60.12 60.65 56.94
MATH 13.00 14.36 14.60 14.76
MiLoRA GSM8K 62.55 63.53 63.99 61.56
MATH 14.12 15.06 16.14 16.16
B.2 More results of the relationship between $\Delta W$ and $W$
Table 11: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs. The weight matrices are taken from the MLP down projection of the 16th layer of LLaMA2-7B.
$\left\|{\bm{W}}\right\|_{F}=123.0$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
$\left\|{\bm{U}}^{\top}{\bm{W}}{\bm{V}}\right\|_{F}=$ 33.34 1.17 1.32 6.98 1.29
$\left\|{\bm{U}}^{\top}\Delta{\bm{W}}{\bm{V}}\right\|_{F}=$ 77.02 74.34 56.61
Amplification factor - - 58.35 10.65 43.88
We present additional results about the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ in Table 11, in which the weight matrices are taken from the MLP down projection of the 16th layer of LLaMA2-7B.
B.3 Math reasoning evaluation results for LLaMA2-7B under PiSSA configuration
Table 12 presents math reasoning evaluation results for LLaMA2-7B finetuned for 10 epoches under PiSSA configuration. The best checkpoint is selected on a validation set.
Table 12: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH finetuned 10 epoches under PiSSA configuration.
Method GSM8K MATH Avg.
LoRA 56.71 10.28 33.50
PiSSA 55.72 10.02 32.87
MiLoRA 55.57 9.08 32.33
Appendix C Limitations and Broader Impacts
Limitations
Due to limited computational resources, following previous work, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks. We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
Broader Impacts
The MiLoRA method enhances model performance with lower training costs. Teams or users with limited computational resources can finetune large models using MiLoRA, promoting the broader application of large models across diverse groups. However, the PEFT methods could potentially be exploited to finetune models for malicious purposes. It is essential to investigate rules and strategies to prevent the misuse of PEFT methods.

### v2

@@
 Abstract
-Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computation and memory cost.
-Previous LoRA-based approaches initialize the low-rank matrices with gaussian distribution and zero values, while keeping the original weight matrices frozen.
-However, the trainable model parameters optimized in an unguided subspace might have interference with the well-learned subspace of the pretrained weight matrix.
-In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principle singular components frozen.
-It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principle matrix contains important knowledge.
-The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principle matrix, thus the pretrained knowledge is expected to be well preserved.
-During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the finetuning dataset.
-Extensive experiments on commonsense reasoning, math reasoning and instruction following benchmarks present the superior performance of our method.
+Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computational and memory cost.
+Previous LoRA-based approaches initialize the low-rank matrices with Gaussian distribution and zero values while keeping the original weight matrices frozen.
+However, the trainable model parameters optimized in an unguided subspace might interfere with the well-learned subspace of the pretrained weight matrices.
+In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principal singular components frozen.
+It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principal matrix contains important knowledge.
+The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principal matrix, thus the pretrained knowledge is expected to be well preserved.
+During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the labeled dataset.
+Extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks present the superior performance of our method.
+MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning
+Hanqing Wang1,  Yixia Li2,  Shuo Wang3,  Guanhua Chen2,  Yun Chen1 1Shanghai University of Finance and Economics 2Southern University of Science and Technology   3Tsinghua University
 1 Introduction
-Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024) and question answering (Ivison et al., 2023).
+Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024a) and question answering (Ivison et al., 2023).
 These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks.
 Fully finetuning is commonly employed to unlock the complete potential of LLMs, however, optimizing all model parameters necessitates substantial and restrictive computational resources (Touvron et al., 2023; Jiang et al., 2023), which hampers the utilization of LLMs across diverse scenarios.
-Parameter-efficient finetuning (Hu et al., 2021, 2023b; Liu et al., 2022; Zhang et al., 2024; Liu et al., 2024a, PEFT) aims at reducing the computational and GPU memory cost for finetuning of pretrained models.
+Parameter-efficient finetuning (Hu et al., 2021, 2023; Liu et al., 2022; Zhang et al., 2024; Liu et al., 2024b, PEFT) aims at reducing the computational and GPU memory cost for finetuning of pretrained models.
 Low-rank adaptation (Hu et al., 2021, LoRA) is one of the most widely used PEFT methods for LLM finetuning.
@@
 For each selected weight matrix, it only updates two low-rank matrices while keeping the pretrained weight frozen.
-During inference, the low-rank matrices are merged into pretrained linear weights, thus no additional computation or memory cost is introduced.
-Recently, researchers explore different LoRA-based variants for efficient LLM finetuning (Zhang et al., 2023; Pan et al., 2024; Kopiczko et al., 2024; Liu et al., 2024a; Meng et al., 2024).
-However, most existing LoRA-based works randomly initialize the low-rank matrices and optimize the trainable parameters in an unguided subspace.
-We suspect this strategy may override important pretrained features, thus influencing the performance of low-rank adaptation methods (Dou et al., 2024).
+During inference, the low-rank matrices are merged into the pretrained linear weights, thus no additional computational or memory cost is introduced.
+Recently, researchers have explored different LoRA-based variants for efficient LLM finetuning (Zhang et al., 2023; Pan et al., 2024; Kopiczko et al., 2024; Liu et al., 2024b; Meng et al., 2024).
+However, most existing LoRA-based methods randomly initialize the low-rank matrices and optimize the trainable parameters in an unguided subspace.
+We suspect this strategy may override important pretrained features, thus degrading the performance of low-rank adaptation methods (Dou et al., 2024).
 In this paper, we propose Minor singular component based Low Rank Adaptation (MiLoRA) for efficient LLM finetuning.
-MiLoRA has a similar framework as vanilla LoRA, but employs a different initialization schedule.
-Specifically, a weight matrix ${\bm{W}}$ is decomposed with singular value decomposition (SVD) algorithm.
-Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principle matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values.
-We argue that the principle matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information.
-It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that using principle low-rank approximation matrices can achieve comparable or even better performance to full parameters.
-Motivated by these observations, we keep the principle matrix ${\bm{W}}_{p}$ frozen and adapt the minor singular components during finetuning.
+MiLoRA has a similar framework as LoRA but employs a different initialization schedule.
+Specifically, a weight matrix ${\bm{W}}$ is decomposed with the singular value decomposition (SVD) algorithm.
+Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principal matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values.
+We argue that the principal matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information.
+It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that principal low-rank approximation can achieve comparable or even better performance to full parameters.
+Motivated by these observations, we freeze the principal matrix ${\bm{W}}_{p}$ and adapt the minor singular components during finetuning.
 The low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ in LoRA framework are initialized with the minor matrix ${\bm{W}}_{m}$.
-Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principle matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge.
+Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principal matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge.
 To maintain the capability of the pretrained model at the start of finetuning, vanilla LoRA explicitly initializes ${\bm{B}}$ with zeros.
-In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals to the frozen principle part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$.
-We conduct extensive experiments on commonsense reasoning, math reasoning and instruction following benchmarks.
-The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+1.92 on LLaMA2-7B), and instruction following (+1.4 on LLaMA2-7B).
+In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals the frozen principal part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$.
+We conduct extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks.
+The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+2.0 on LLaMA2-7B), instruction following (+1.4 on LLaMA2-7B) and visual instruction following (+1.4 on LLaVA1.5-7B).
 [^1] Our code and model will be made public.
 2 Preliminaries
-Singualr Value Decomposition
-Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition (SVD) is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$.
+Singular Value Decomposition
+Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$.
 The columns of ${\bm{U}}$ are the left singular vectors, and the columns of ${\bm{V}}$ are the right singular vectors.
 The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order.
-Without loss of generality, we suppose $m\leq n$ to simplify the notation.
-The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
+Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
 ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}=\sum_{i=1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$ (1)
@@
 The LoRA method only modifies the linear matrices in the Transformer model.
-The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with fully finetuning.
-However, vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized.
-This might potentially detract from the pretrained knowledge encoded in the principal weight matrix.
+The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with full finetuning.
+However, the vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized.
+This might potentially detract from the pretrained knowledge encoded in the pretrained weight matrix.
 Figure 1: (a) MiLoRA method splits the pretrained linear weight matrix into two parts, the minor singular component (${\bm{U}}_{m},{\bm{V}}_{m},{\bm{\Sigma}}_{m}$) is used to initialize the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ in the LoRA framework.
-(b) During finetuning, only the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ are updated while the principle matrix ${\bm{W}}_{p}$ is frozen, which is similar as the practice of LoRA method.
+(b) During finetuning, only the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ are updated while the principal matrix ${\bm{W}}_{p}$ is frozen, which is similar to the practice of LoRA method.
 3 Methodology
 The proposed MiLoRA is a simple yet effective PEFT approach that selects the minor singular components of the weight matrices for optimization.
-As observed in LASER (Sharma et al., 2024), the minor singular components of weight matrices contain noisy or long-tail information, while the principle singular components are responsible for important features across tasks.
+As observed in LASER (Sharma et al., 2024), the minor singular components of weight matrices contain noisy or long-tail information, while the principal singular components are responsible for important features across tasks.
 Therefore, MiLoRA is designed to effectively learn from the finetuning dataset by adapting minor singular components, while maintaining the knowledge encoded in the pretrained model.
@@
 Specifically, suppose $m\leq n$, MiLoRA splits each linear weight matrix into two matrices according to the corresponding singular values: principal matrix ${\bm{W}}_{p}$ and minor matrix ${\bm{W}}_{m}$ (see Figure 1):
-${\bm{W}}={\bm{W}}_{p}+{\bm{W}}_{m}=\sum_{i=1}^{m-r}\sigma_{i}u_{i}v_{i}^{\top}+\sum_{i=m-r+1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$ (3)
+${\bm{W}}$ $={\bm{W}}_{p}+{\bm{W}}_{m}$ (3)
+$=\sum_{i=1}^{m-r}\sigma_{i}u_{i}v_{i}^{\top}+\sum_{i=m-r+1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$
 where the singular values are in descending order ($\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}\geq 0$), and the hyperparameter $r$ is the number of minor singular values in the ${\bm{W}}_{m}$ matrix.
 The decomposition can be also reformulated in a matrix form.
-The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principle and minor singular values, respectively.
-The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated in a similar way.
+The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principal and minor singular values, respectively.
+The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated similarly.
 The SVD of ${\bm{W}}$ can be expressed as:
-${\bm{W}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}={\bm{U}}_{p}{\bm{\Sigma}}_{p}{\bm{V}}_{p}^{\top}+{\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}={\bm{W}}_{p}+{\bm{W}}_{m}$ (4)
+${\bm{W}}$ $={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}={\bm{U}}_{p}{\bm{\Sigma}}_{p}{\bm{V}}_{p}^{\top}+{\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}$ (4)
+$={\bm{W}}_{p}+{\bm{W}}_{m}$
 During the finetuning process, instead of freezing the entire pretrained weight matrix, we just keep the principal singular components ${\bm{W}}_{p}$ fixed to preserve the pretrained knowledge.
 The minor matrix ${\bm{W}}_{m}$ is used to initialize low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ (see Figure 1):
-${\bm{W}}_{m}={\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}=({\bm{U}}_{m}\sqrt{{\bm{\Sigma}}_{m}})(\sqrt{{\bm{\Sigma}}_{m}}{\bm{V}}_{m}^{\top})={\bm{B}}_{m}{\bm{A}}_{m}$ (5)
-This strategy has two benefits: 1) It encourages the model to learn in the less-optimized subspace spanned by the minor singular vectors, thus reducing the interference with the well-learned pretrained knowledge encoded in the principle singualr components.
+${\bm{W}}_{m}$ $={\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}$ (5)
+$=({\bm{U}}_{m}\sqrt{{\bm{\Sigma}}_{m}})(\sqrt{{\bm{\Sigma}}_{m}}{\bm{V}}_{m}^{\top})={\bm{B}}_{m}{\bm{A}}_{m},$
+This strategy has two benefits: 1) It encourages the model to learn in the less-optimized subspace spanned by the minor singular vectors, thus reducing the interference with the well-learned pretrained knowledge encoded in the principal singular components.
 2) Unlike vanilla LoRA which requires tuning the scaling factor $\alpha$ and the initialization hyperparameters, ours does not require any hyperparameter except the rank $r$.
 To make our method simple and training-efficient, we do not incorporate the regularization loss which imposes strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$.
-Our experiment results show that our scheme works quite well in balancing learning from finetuning dataset and preverving pretrained knowledge.
+Our experiment results show that our scheme works quite well in balancing learning from the finetuning dataset and preserving pretrained knowledge.
 4 Experiments
-To evaluate our method against other PEFTs, we conduct experiments across three diverse tasks covering more than 10 datasets.
-Our objective is to illustrate the effectiveness of our method across various scenarios.
-We benchmark against LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024).
-  • LoRA (Hu et al., 2021) reparameterizes the weight update $\Delta{\bm{W}}$ with two trainable low-rank matrices ${\bm{A}}$ and ${\bm{B}}$, while freezing the pretrained weight ${\bm{W}}$.
-They use a random Gaussian initialization for A and zero for B, so $\Delta{\bm{W}}={\bm{B}}{\bm{A}}$ is zero at the beginning of training.
-  • PiSSA (Meng et al., 2024) shares the same framework as LoRA, but it employs a more sophisticated initialization approach.
-Specifically, they initialize ${\bm{A}}$ and ${\bm{B}}$ with principal singular values and singular vectors of the pre-trained weight ${\bm{W}}$.
-Given that the principal singular values and vectors capture the essence of a matrix, PiSSA is hoped to better approximate the outcomes of full-parameter finetuning by changing the essential parts while freezing the “noisy” parts.
-The MiLoRA method optimizes the low-rank matrices within a less-optimized subspace, which can be combined with other LoRA-based PEFT methods (Liu et al., 2024a; Zhang et al., 2023; Pan et al., 2024).
-We leave the exploration of combination MiLoRA with LoRA variants as future work.
-4.1 Commonsense reasoning
-Setup
-We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023a).
-Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018).
-The task is formulated as a multiple-choice problem.
-The best checkpoint is chosen by their validation set loss.
-We report accuracy (%) for all datasets.
-We use the implementation of LLM-Adapters (Hu et al., 2023b) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters.
-Hyperparameter
-We use the same hyperparameter configurations as Hu et al. (2023a) without tuning for all methods. Details can be found in Table 8 of Appendix A.2. We denote this hyperparameter setup as LLM-Adapters.
-Results
-Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks.
-Performance results of baseline methods are taken from Liu et al.
-(2024a).
-Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
 Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
-ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
+ChatGPT -- 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
 LLaMA2-7B LoRA 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
@@
 MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
-We report results in Table 1.
-We also include the ChatGPT baseline reported in Liu et al.
-(2024a) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022).
-MiLoRA consistently surpasses all baseline methods across both LLaMA2-7B and LLaMA3-8B models.
-On LLaMA2-7B, MiLoRA outperforms LoRA and PiSSA by 1.6 and 5.4 average accuracy scores.
-On LLaMA3-8B, MiLoRA exceeds LoRA and PiSSA by 1.1 and 6.5 average accuracy scores.
-These outcomes suggest the effectiveness of MiLoRA as a PEFT method.
-4.2 Math reasoning
-Setup
-We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH(Hendrycks et al., 2021).
-We use test set of GSM8K and MATH for evaluation and report results on the last checkpoint.
-For each test set, we report the exact match ratio against the ground truth.
-We use the implementation of PiSSA(Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA.
-Hyperparameter
-As MetaMathQA has a large training dataset, we set the rank to 64.
-For all other hyperparameters, we follow the llm-adapter hyperparameter setup shown in Table 8 of Appendix A.2.
-We limit all samples to a maximum length of 2048 tokens.
-Results
-Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH. Results are cited from Yu et al. (2023).
+Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks.
+Results are cited from Liu et al.
+(2024b).
+Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
 Method GSM8K MATH Avg.
 Full FT 66.5 19.8 43.2
-LoRA 60.58 16.88 38.73
-PiSSA 58.23 15.84 37.04
-MiLoRA 63.53 17.76 40.65
-We report results in Table 2.
-We also include the full finetuning results from (Yu et al., 2023) as a reference.
-We can see that MiLoRA achieves the best performance on LLaMA2-7B.
-However, it still underperforms full finetuning by an average of 2.55 EM score.
-This suggests that PEFT methods still have room for improvement to fully match the performance of full finetuning.
-4.3 Instruction-following
-Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Performance results of baseline methods are taken from Ivison et al. (2023).
+LoRA 60.6 16.9 38.7
+PiSSA 58.2 15.8 37.0
+MiLoRA 63.5 17.8 40.7
+Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH. Results are cited from Yu et al. (2023).
+To evaluate MiLoRA against other PEFTs, we conduct experiments on both NLP and multimodal tasks covering 18 datasets.
+We benchmark against LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024).
+All of our experiments are conducted on 8 NVIDIA L40 GPUs unless otherwise specified.
+  • LoRA (Hu et al., 2021) reparameterizes the weight update $\Delta{\bm{W}}$ with two trainable low-rank matrices ${\bm{A}}$ and ${\bm{B}}$, while freezing the pretrained weight ${\bm{W}}$.
+They use a random Gaussian initialization for A and zero for B, so $\Delta{\bm{W}}={\bm{B}}{\bm{A}}$ is zero at the beginning of training.
+  • PiSSA (Meng et al., 2024) shares the same framework as LoRA, but employs a more sophisticated initialization approach.
+They initialize ${\bm{A}}$ and ${\bm{B}}$ with principal singular values and singular vectors of the pre-trained weight ${\bm{W}}$.
+Given that the principal components capture the essence of a matrix, PiSSA is expected to better approximate full finetuning by changing the essential parts while freezing the “noisy” parts.
+4.1 Experiments on Large Language Model
+Models and Datasets
+We evaluate and compare MiLoRA with baselines on three different types of LLM downstream tasks.
+1.
+Commonsense reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023).
+Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018).
+The task is formulated as a multiple-choice problem.
+We report accuracy (%) for all datasets on the best checkpoint chosen by the validation set loss.
+2.
+Math reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
+We use test sets of GSM8K and MATH for evaluation and report results on the last checkpoint. For each test set, we report the exact match ratio against the ground truth.
+3.
+Instruction following: We follow the experiment setup in Wu et al.
+(2024a) and Wu et al.
+(2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023).
+Following previous works, we use Alpaca-Eval v1.0 (Li et al., 2023a) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024).
+Implementation Details
+We use the same hyperparameter configurations as Hu et al.
+(2023) without tuning for all methods.
+Details can be found in Appendix A.2.
+We denote this hyperparameter setup as LLM-Adapters.
+By default, we use a rank of 32.
+As math reasoning has a large training dataset, we set the rank to 64.
 Method Win-rate
@@
 GPT-3.5-turbo-0613 91.8
-LLaMA2 Chat 7B 87.3
-LLaMA2 7B + LoRA 85.9
-LLaMA2 7B + PiSSA 85.6
-LLaMA2 7B + MiLoRA 87.3
-Setup
-We follow the experiment setup in Wu et al.
-(2024a) and Wu et al.
-(2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023).
-Following previous works, we use Alpaca-Eval v1.0  (Li et al., 2023) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024).
-We implement our method using open-instruct [^4] https://github.com/allenai/open-instruct.
-Hyperparameter
-We use the LLM-Adapters hyperparameter setup for all PEFT methods.
-We truncate samples to a maximum of 768 tokens.
-During evaluation, we use the last checkpoint with a greedy decoding strategy and a maximum token number of 2048.
-To avoid the issue of repetition, we set the repetition penalty to 1.1 and set the no repeat n-gram size to 5, following Wu et al.
-(2024a) and Wu et al.
+LLaMA2-Chat-7B 87.3
+LLaMA2-7B + LoRA 85.9
+LLaMA2-7B + PiSSA 85.6
+LLaMA2-7B + MiLoRA 87.3
+Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Results are cited from Ivison et al. (2023).
+Results
+Method VQAv2 GQA VisWiz SQA VQAT POPE MMBench Avg.
+Full FT 78.5 61.9 50.0 66.8 58.2 85.9 64.3 66.5
+LoRA 79.1 62.9 47.8 68.4 58.2 86.4 66.1 66.9
+PiSSA 77.5 60.6 39.2 67.3 54.3 87.0 61.4 63.9
+MiLoRA 79.2 62.1 53.3 70.6 58.7 87.9 66.1 68.3
+Table 4: Visual instruction tuning evaluation results for LLaVA1.5-7B on seven vision-language tasks. We directly use checkpoints from Liu et al. (2024a) to reproduce the results of baseline methods.
+We report results in Table 1, Table 2 and 3 for commonsense reasoning, math reasoning and instruction following, respectively.
+For commonsense reasoning, we also include the ChatGPT baseline reported in Liu et al.
+(2024b) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022).
+For math reasoning, we add the full finetuning results from Yu et al.
+(2023) as a reference.
+For instruction following, we include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2-Chat-7B from Ivison et al.
+(2023) as a reference.
+As can be seen, MiLoRA consistently surpasses all baseline methods across different datasets and LLMs, indicating that MiLoRA is a highly effective finetuning method.
+Specifically, for commonsense reasoning, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 5.4 accuracy points on LLaMA2-7B, and by 1.1 and 6.5 points on LLaMA3-8B.
+For math reasoning, MiLoRA improves over LoRA and PiSSA by 2.0 and 3.7 averaged EM points on LLaMA2-7B, respectively.
+However, it still underperforms full finetuning by an average of 2.5 EM score.
+This suggests that the PEFT methods still have room for improvement to fully match the performance of full finetuning.
+For instruction following, MiLoRA achieves the best results, surpassing LoRA and PiSSA by 1.4 and 2.7 in win-rate scores, respectively.
+4.2 Experiments on Vision-Language Model
+Models and Datasets
+To further examine if MiLoRA can remain competitive on multimodal finetuning tasks, we further conduct visual instruction tuning tasks on LLaVA1.5-7B Liu et al.
+(2024a), which consists of an LLM, Vicuna-1.5-7B Peng et al.
+(2023) and a vision encoder, CLIP ViT-L/336px Radford et al.
+(2021).
+The visual instruction tuning is performed in a multitask setting, including VQA Hudson and Manning (2019); Marino et al.
+(2019); Schwenk et al.
+(2022), OCR Mishra et al.
+(2019); Sidorov et al.
+(2020), region-level VQA Krishna et al.
+(2017); Mao et al.
+(2016), visual conversation Liu et al.
+(2024a), and language conversation data.
+The finetuned models are evaluated on seven different multimodal benchmarks: VQA-v2 Goyal et al.
+(2017), GQA Hudson and Manning (2019), VizWiz Gurari et al.
+(2018), SQA Lu et al.
+(2024), VQAT Singh et al.
+(2019), POPE Li et al.
+(2023b), and MMBench Liu et al.
+(2024c).
+Implementation Details
+We follow the settings of Liu et al.
+(2024a) to construct the training data and prompt template.
+For the hyperparameter configuration, we follow the LoRA configuration provided by Liu et al.
+(2024a) without tuning for all PEFT methods for a fair comparison.
+Details can be found in Appendix A.2.
+Results
+Table 4 shows the experiment results.
+As can be seen, MiLoRA achieves the best results, outperforming LoRA and PiSSA by 1.4 and 4.4 average accuracy points, respectively.
+We also note that in this setup, fully fine-tuning is less effective than LoRA.
+As a result, methods like PiSSA, which aim to closely approximate fully fine-tuning, may lose their advantage.
+In contrast, MiLoRA improves performance by retaining more of the pre-trained knowledge, thereby circumventing this issue.
+5 Understanding MiLoRA
+In this section, we conduct experiments to further understand MiLoRA. By default, we use the finetuned LLaMA2-7B model on math reasoning from our main experiments for analyses.
+5.1 Comparison with More PEFTs
+To further demonstrate the effectiveness of MiLoRA, we compare MiLoRA with additional popular baselines, such as rsLoRA Kalajdzievski (2023), LoRA+ Hayou et al.
+(2024), DoRA Liu et al.
+(2024b), AdaLoRA Zhang et al.
+(2023), LoRA-GA Wang et al.
 (2024b).
-Results
-Table 3 presents the win-rate of various PEFT methods.
-We include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2 Chat 7B as a reference, which have win-rate of 91.2%, 91.8%, and 87.3% respectively.
-Our method, LLaMA2-7B + MiLoRA outperforms the other two PEFT methods and matches LLaMA2 Chat 7B’s performance, indicating that MiLoRA is a highly effective finetuning method.
-5 Understanding MiLoRA
-In this section, we conduct experiments to further understand MiLoRA.
-For all analysis experiments that do not involve finetuning, we use the LLaMA2-7B model finetuned on MetaMathQA in our main experiments (see Section 4.2).
-For any experiments that require finetuning except for Section 5.5, we finetune the LLaMA2-7B model on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512.
-This configuration is chosen to minimize computational costs.
-5.1 How do different rank settings influence MiLoRA?
-Figure 2: Average score of finetuned LLaMA2-7B model on math reasoning across different ranks.
-In Figure 2, we explore the impact of various rank configurations by adjusting $r$ within the set $\{16,32,64,128\}$ and assessing the average accuracy of the finetuned LLaMA2-7B model on math reasoning tasks.
-Detailed numbers are presented in Table 10 of Appendix B.1.
-As can be seen, MiLoRA with rank 64 achieves the best performance across all methods and ranks.
-MiLoRA consistently surpasses PiSSA across all rank configures.
-However, when compared with LoRA, MiLoRA performs better only at large ranks.
-This is likely due to the fact that MiLoRA makes fewer changes to the pre-trained parameters compared to LoRA (see detailed discussion in Section 5.3 and 5.4).
-As a result, when the rank is relatively low, MiLoRA struggles to fully digest the finetuning dataset.
+We follow the math and code experiment settings in Wang et al.
+(2024b).
+Specifically, we finetune LLaMA2-7B with MiLoRA on MetaMathQA 100K and Code-Feedback 100K for one epoch using rank $8$ and compare MiLoRA with baseline results reported in Wang et al.
+(2024b).
+Table 5 shows the results.
+We report the training cost of all methods relative to LoRA, which requires 3.58 hours of training on the math task using a single NVIDIA-L40 GPU.
+As can be seen, MiLoRA consistently outperforms all baselines, improving the best PEFT baseline LoRA-GA by 1.1 and 4.6 points on GSM8K and Human-eval, respectively.
+The training efficiency of MiLoRA is also very high.
+The time for SVD decomposition in MiLoRA is less than six minutes and negligible when compared with the total training time.
+Therefore, the training time required of MiLoRA is consistent with that of LoRA, while DoRA and AdaLoRA require additional training time overhead.
+Method GSM8K Human-eval Training Cost
+LoRA 42.1 14.8 $\times$ 1.00
+PiSSA 44.5 16.0 $\times$ 1.00
+rsLoRA 45.6 16.0 $\times$ 1.00
+LoRA+ 52.1 18.2 $\times$ 1.00
+DoRA 53.1 19.8 $\times$ 1.58
+AdaLoRA 50.7 17.8 $\times$ 1.07
+LoRA-GA 53.6 19.8 $\times$ 1.00
+MiLoRA 54.7 24.4 $\times$ 1.00
+Table 5: Comparison with more baselines on math and code tasks. Baseline GSM8K and Human-eval results are cited from Wang et al. (2024b).
 5.2 Is Minor Singular Component Adaptation Important?
 To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model.
-Specifically, we initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and evaluate the results on the GSM8K and MATH datasets.
-Table 4: Performance of MiLoRA when initializing with principal, random sampled, and minor singular components.
-Method Principle Random Minor
-GSM8K 60.65 63.23 63.99
-MATH 14.60 15.54 16.14
-The results, shown in Table 4, demonstrate that initializing with the minor singular components achieves the best performance across both evaluation datasets.
+Specifically, we finetune LLaMA2-7B on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512 to save computation.
+We initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and report the evaluation results in Table 6.
+As can be seen, initializing with the minor singular components achieves the best performance across both datasets.
 This underscores the importance of adapting the minor singular components during the finetuning process.
 Concurrent with our work, Meng et al.
-(2024) propose PiSSA, a similar approach to MiLoRA but adapts the principle singular components.
+(2024) propose PiSSA, a similar approach to MiLoRA but adapts the principal singular components.
 We will provide a detailed comparison between MiLoRA and PiSSA in Section 5.5.
+Method principal Random Minor
+GSM8K 60.7 63.2 64.0
+MATH 14.6 15.5 16.1
+Table 6: Performance of MiLoRA when initializing with principal, random sampled, and minor singular components.
 5.3 How does the Matrix Update $\Delta W$ Compare to $W$?
@@
 We make two key observations.
-First, $\Delta{\bm{W}}$ of LoRA and MiLoRA has a similar correlation with ${\bm{W}}$, both slightly stronger compared to a random matrix, indicating that they amplify directions that are not emphasized in ${\bm{W}}$.
+First, $\Delta{\bm{W}}$ of LoRA and MiLoRA have a similar correlation with ${\bm{W}}$, both slightly stronger than a random matrix, indicating that they amplify directions not emphasized in ${\bm{W}}$.
 In contrast, $\Delta{\bm{W}}$ of PiSSA shows a strong correlation with ${\bm{W}}$.
@@
 Compared to LoRA, we find the amplification factor of MiLoRA is significantly smaller.
-This indicates that MiLoRA has a reduced impact on features already present in ${\bm{W}}$.
-See Table 11 of Appendix B.2 for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
-Table 5: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs.
-The weight matrices are taken from the query projection of the 16th layer of LLaMA2-7B.
+This indicates that MiLoRA has a reduced impact on the features which already present in ${\bm{W}}$.
+See Table 13 of Appendix B for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
 $\left\|{\bm{W}}\right\|_{F}=97.36$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
@@
 Amplification factor - - 37.46 3.18 24.17
+Table 7: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs.
+The weight matrices are taken from the query projection of the 16th layer of LLaMA2-7B.
+Figure 2: Left: The subspace similarity averaged by all modules in a layer. Right: The subspace similarity averaged by all layers for a specific module type.
 To further analyze the characteristics of $\Delta{\bm{W}}$ for different methods, we measure the normalized subspace similarity (Hu et al., 2021) between the subspaces spanned by top $r$ singular vectors of $\Delta{\bm{W}}$ and the subspace spanned by top $r$ singular vectors, bottom $r$ singular vectors, and a random $r$ singular vectors of ${\bm{W}}$.
@@
 The right figures further show that MiLoRA places greater emphasis on features associated with the bottom singular vectors of the MLP layers.
-Figure 3: Left: The subspace similarity averaged by all modules in a layer. Right: The subspace similarity averaged by all layers for a specific module type.
-5.4 Does MiLoRA forget less than baselines?
-Table 6: Forgetting loss of various PEFT methods.
+5.4 MiLoRA Forgets Less than Baselines
 Method LoRA PiSSA MiLoRA
 Forgetting loss 3.24 6.07 2.54
+Table 8: Forgetting loss of various PEFT methods.
 A hypothesis explanation for why MiLoRA outperforms baselines is that pre-trained knowledge is more retained.
@@
 This is the usual next token prediction loss used when training LLMs, except that the target next token is replaced by the distribution predicted by the pre-trained base model.
-We evaluate the forgetting metric on LLaMA2-7B model finetuned in Section 4.2 using the WikiText-103 test dataset following (Kalajdzievski, 2024).
+We evaluate the forgetting metric on the LLaMA2-7B model finetuned in math reasoning using the WikiText-103 test dataset following Kalajdzievski (2024).
 As shown in Table 8, MiLoRA exhibits the lowest forgetting loss, which is consistent with our hypothesis that MiLoRA makes the least modification to pre-trained knowledge.
@@
 5.5 Comparison Between MiLoRA and PiSSA
+Hyp. Setup PEFT GSM8K MATH Avg.
+PiSSA LoRA 41.5 5.8 23.6
+PiSSA 51.3 7.6 29.4
+MiLoRA 40.0 5.2 22.6
+LLM-Adapters LoRA 56.6 10.8 33.7
+PiSSA 51.3 10.4 30.8
+MiLoRA 58.6 11.6 35.1
+Table 9: Math reasoning evaluation results for LLaMA2-7B with PiSSA hyperparameters and our (LLM-Adapters) hyperparameters.
 Concurrent with our research, Meng et al.
-(2024) have recently proposed a low-rank adaptation method called PiSSA.
-At first glance, PiSSA may appear to be quite similar to MiLoRA but adapts the principle singular components.
-However, there are fundamental differences between PiSSA and MiLoRA that are worth highlighting:
+(2024) have recently introduced a low-rank adaptation method called PiSSA.
+PiSSA shares a similar framework as MiLoRA, but adapts the principal singular components.
+We argue that PiSSA and MiLoRA are fundamentally different.
 Motivation: PiSSA is designed to approximate full finetuning by adapting the principal singular components, which are believed to capture the essence of the weight matrices.
@@
 Second, we find that all methods perform better in our hyperparameter setup, suggesting that our hyperparameter configuration is more effective.
-Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.39 and 4.27 average scores, respectively.
+PiSSA configuration of hyperparameters has a combination of a small learning rate (2e-5) and a large batch size (128).
+This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023; Biderman et al., 2024).
+Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.4 and 4.3 average scores, respectively.
 In summary, when using an appropriate hyperparameter configuration, our proposed MiLoRA method achieves markedly superior performance compared to PiSSA.
-We further investigate why the hyperparameter configuration in PiSSA paper is suboptimal.
-This configuration, as shown Table 9 of Appendix A.3, is from the full finetuning hyperparameters of Alpaca (Taori et al., 2023).
-As can be seen, it has a combination of small learning rate (2e-5) and large batch size (128).
-This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023).
-This problem becomes more severe as PiSSA paper only trains for a single epoch for all experiments.
-Figure 4: Validation loss curve of different PEFTs under PiSSA hyperparameter configurations. LoRA and MiLoRA are under-fitting when training with 1 epoch (781 steps).
-To validate our assumption, we extract another 120 samples from MetaMathQA to serve as a validation set and then train LoRA, PiSSA, and MiLoRA under the same hyperparameter configuration except for 10 epochs each and recorded their respective validation loss curves.
-From the results shown in Figure 4, we make two key observations.
-First, LoRA and MiLoRA are much under-fitting when training with only 1 epoch (781 steps).
-However, with sufficient training, their validation loss can converge to the same level as that of PiSSA.
-Second, PiSSA converges faster than the other two methods but quickly overfits.
-This might be attributed to that PiSSA directly finetunes the essence of weight via the top singular vectors and values.
-We evaluate the best checkpoint of each methods selected by the validation set, and find that LoRA performs best, while the other two perform similarly (see Table 12 in Appendix B.3), which again confirms that LoRA and MiLoRA are observed worse than PiSSA in the PiSSA paper because they are underfitting.
-Table 7: Math reasoning evaluation results for LLaMA2-7B with PiSSA hyperparameters and our (LLM-Adapters) hyperparameters.
-Hyp. Setup PEFT GSM8K MATH Avg.
-PiSSA LoRA 41.47 5.76 23.6
-PiSSA 51.25 7.58 29.4
-MiLoRA 40.03 5.18 22.6
-LLM-Adapters LoRA 56.56 10.84 33.70
-PiSSA 51.25 10.38 30.82
-MiLoRA 58.61 11.56 35.09
 6 Related Work
 Parameter-Efficient Finetuning of LLMs.
-After pretraining, large language models are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023).
-Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of fully finetuning.
-Previous PEFT methods can be grouped into three lines.
-The first group is adapter-based, which inserts additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b).
-However, adapter-based methods introduce additional computation costs during inference.
-The second category is LoRA-based (Hu et al., 2021; Liu et al., 2024a; Zhang et al., 2023), which models the changes of selected linear weights as low-rank matrices.
+After pretraining, LLMs are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023).
+Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of full finetuning.
+Previous PEFT methods can be grouped into three lines: adapter-, LoRA- and prompt-based methods.
+Adapter-based methods insert additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b).
+However, they introduce additional computational costs during inference.
+LoRA-based methods (Hu et al., 2021; Liu et al., 2024b; Zhang et al., 2023) model the changes of selected linear weights as low-rank matrices.
 During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen.
-The third group is prompt-based (Lester et al., 2021; Li & Liang, 2021; Liu et al., 2022) which adds additional soft prompts to the input tokens.
+Prompt-based methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022) add additional soft prompts to the input tokens.
 During training, only soft prompts are updated while the pretrained model parameters are fixed.
-Among these PEFT methods, the LoRA-based approaches are most widely used for LLM finetuning as it does not introduce computation overhead during inference while approximating the performance of fully finetuning.
+Among these PEFT methods, the LoRA-based approaches are widely used for LLM fine-tuning because they are easy to implement and do not introduce computational overhead during inference.
 LoRA and Its Variants
@@
 With this lightweight decomposition, LoRA reduces storage and task-switching overhead by sharing the pretrained models across multiple tasks.
-Theoretically, the linear design of LoRA enables seamless merging of trainable matrices with frozen weights, thereby avoiding extra inference latency.
-Since then, researchers explore and propose different LoRA variants for PEFT.
-AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024b), adaptively determine the rank of LoRA module in each weight matrix according to the importance score.
-The pruned ranks are compensated on the more important positions.
+Since then, researchers have explored and proposed different LoRA variants for PEFT.
+AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024d), adaptively determine the rank of LoRA module in each weight matrix according to the importance score.
 The rsLoRA (Kalajdzievski, 2023) modifies the LoRA with the appropriate scaling factor to improve the performance of large ranks.
-The DoRA (Liu et al., 2024a) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance.
-VeRA (Kopiczko et al., 2024) reduces the number of trainable parameters compared to LoRA by sharing a pair of low-rank matrices across all layers.
-Concurrent with our work, PISSA (Meng et al., 2024) proposes to only update the principle singular components.
-However, their conclusion is empirically problematic due to the sub-optimal hyperparameters.
+The DoRA (Liu et al., 2024b) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance.
+Concurrent with our work, PISSA (Meng et al., 2024) and LoRA-GA Wang et al.
+(2024b) propose to better approximate full finetuning by only updating the principal singular components or aligning the gradients of low-rank updates to that of full finetuning.
 7 Conclusion
@@
 MiLoRA effectively learns on finetuning tasks while better preserving the pretrained knowledge by adapting the minor singular components of pretrained weight matrices.
-We investigate the effectiveness of MiLoRA on a wide range of LLM evaluation benchmarks, including commonsense reasoning, math reasoning and instruction-following.
+We investigate the effectiveness of MiLoRA on a wide range of LLM evaluation benchmarks, including commonsense reasoning, math reasoning and instruction-following, and vision-language model evaluation benchmarks.
 Experiment results demonstrate that MiLoRA consistently outperforms LoRA and PiSSA without sacrificing training or inference efficiency.
 We hope that our work will inspire future research on parameter-efficient finetuning of LLMs.
+Limitations
+Due to limited computational resources, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks following previous work Liu et al.
+(2024b); Wu et al.
+(2024b).
+We examine and compare MiLoRA with baselines using the hyperparameter configurations provided in the previous works Hu et al.
+(2023); Liu et al.
+(2024b) instead of exhaustive hyperparameter search for each task.
+We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
 Appendix A Detailed Experiment Setups
-A.1 Similarity metric between subspace
+A.1 Similarity Metric Between Subspaces
 Following Hu et al.
@@
 Given two matrices $\Delta{\bm{W}}$ and ${\bm{W}}$, we extract $r$ left singular vectors from each to form the subspace matrices for them, which are denoted as $\Delta{\bm{W}}_{s}$ and ${\bm{W}}_{s}$ , and then the subspace similarity is computed using $\phi(\Delta{\bm{W}}_{s},{\bm{W}}_{s})$.
-A.2 Our hyperparameter setup
-Table 8 shows our detailed hyperparameters. This setup follows LLM-adapters (Hu et al., 2023b), therefore we denote it as LLM-adapters setup.
-Table 8: Our hyperparameter configuration on the commonsense reasoning (ComR), math reasoning (MathR) and instruction-following (InsF) tasks.
+A.2 Our Hyperparameter Setup for LLM
+Table 10 shows our detailed hyperparameters. This setup follows LLM-adapters (Hu et al., 2023), therefore we denote it as LLM-adapters setup.
 Hyperparameters ComR MathR InsF
@@
 Placement query, key, value, MLP up, MLP down
-A.3 PiSSA hyperparameter setup
+Table 10: Our hyperparameter configuration on the commonsense reasoning (ComR), math reasoning (MathR) and instruction-following (InsF) tasks.
+Visual instruction tuning hyperparameters
+Rank r 128
+$\alpha$ of LoRA 256
+$\alpha$ of PiSSA/MiLoRA 128
+Dropout 0.05
+Optimizer AdamW
+LR 2e-4
+LR Scheduler cosine
+Batch Size 16
+Warmup Ratio 0.03
+Epochs 1
+Placement query, key, value, output, gate, MLP up, MLP down
+Table 11: Our hyperparameter configurations applied in visual instruction tuning tasks.
+A.3 PiSSA Hyperparameter Setup for LLM
 Table 12 shows the detailed hyperparameters from the PiSSA paper, which we denoted as PiSSA hyperparameter setup.
-Table 9: Hyperparameter configuration from the PiSSA paper.
 PiSSA hyperparameters
@@
 Placement query, key, value, output, gate, MLP up, MLP down
-A.4 Computation resource
-All of our experiments are conducted on 8 NVIDIA L40 GPUs.
+Table 12: The detailed hyperparameter configurations used by PiSSAMeng et al. (2024).
+A.4 More Implementation Details
+We use the implementation of LLM-Adapters (Hu et al., 2023) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters for commonsense reasoning, the implementation of PiSSA (Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA for math reasoning and the implementation of open-instruct [^4] https://github.com/allenai/open-instruct for instruction tuning.
+For hyperparameters of finetuning the vision-language model, we follow the LoRA configuration provided by Liu et al.
+(2024a) without tuning for all PEFT methods for a fair comparison.
+Table 11 shows the details.
 Appendix B More Experiment Results
-B.1 Detailed results on math reasoning across ranks
-Detailed EM scores of MiLoRA and baselines on math reasoning by adjusting rank $r$ within the set $\{16,32,64,128\}$.
-Table 10: Math reasoning evaluation results for LLaMA2-7B across different ranks.
-Method Dataset 16 32 64 128
-LoRA GSM8K 62.93 63.38 63.61 59.59
-MATH 15.52 15.56 15.28 15.28
-PiSSA GSM8K 61.56 60.12 60.65 56.94
-MATH 13.00 14.36 14.60 14.76
-MiLoRA GSM8K 62.55 63.53 63.99 61.56
-MATH 14.12 15.06 16.14 16.16
-B.2 More results of the relationship between $\Delta W$ and $W$
-Table 11: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs.
-The weight matrices are taken from the MLP down projection of the 16th layer of LLaMA2-7B.
 $\left\|{\bm{W}}\right\|_{F}=123.0$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
@@
 Amplification factor - - 58.35 10.65 43.88
+Table 13: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs.
+The weight matrices are taken from the MLP down projection of the 16th layer of LLaMA2-7B.
 We present additional results about the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ in Table 13, in which the weight matrices are taken from the MLP down projection of the 16th layer of LLaMA2-7B.
-B.3 Math reasoning evaluation results for LLaMA2-7B under PiSSA configuration
-Table 12 presents math reasoning evaluation results for LLaMA2-7B finetuned for 10 epoches under PiSSA configuration. The best checkpoint is selected on a validation set.
-Table 12: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH finetuned 10 epoches under PiSSA configuration.
-Method GSM8K MATH Avg.
-LoRA 56.71 10.28 33.50
-PiSSA 55.72 10.02 32.87
-MiLoRA 55.57 9.08 32.33
-Appendix C Limitations and Broader Impacts
-Limitations
-Due to limited computational resources, following previous work, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks.
-We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
-Broader Impacts
+Appendix C Broader Impacts
 The MiLoRA method enhances model performance with lower training costs.

### v3

@@
 Abstract
-Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computational and memory cost.
+Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computational and memory costs.
 Previous LoRA-based approaches initialize the low-rank matrices with Gaussian distribution and zero values while keeping the original weight matrices frozen.
@@
 Extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks present the superior performance of our method.
-MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning
-Hanqing Wang1,  Yixia Li2,  Shuo Wang3,  Guanhua Chen2,  Yun Chen1 1Shanghai University of Finance and Economics 2Southern University of Science and Technology   3Tsinghua University
+MiLoRA: Harnessing Minor Singular Components for
+Parameter-Efficient LLM Finetuning
+Hanqing Wang1,  Yixia Li2,  Shuo Wang3,  Guanhua Chen2thanks: Corresponding Authors.
+,  Yun Chen1,411footnotemark: 1 1Shanghai University of Finance and Economics 2Southern University of Science and Technology,  3Tsinghua University 4MoE Key Laboratory of Interdisciplinary Research of Computation and Economics, Shanghai University of Finance and Economics
 1 Introduction
-Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024a) and question answering (Ivison et al., 2023).
+Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024b) and question answering (Ivison et al., 2023).
 These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks.
@@
 We argue that the principal matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information.
-It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that principal low-rank approximation can achieve comparable or even better performance to full parameters.
+It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024; Li et al., 2024) that principal low-rank approximation can achieve comparable or even better performance to full parameters.
 Motivated by these observations, we freeze the principal matrix ${\bm{W}}_{p}$ and adapt the minor singular components during finetuning.
@@
 We conduct extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks.
-The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+2.0 on LLaMA2-7B), instruction following (+1.4 on LLaMA2-7B) and visual instruction following (+1.4 on LLaVA1.5-7B).
-[^1] Our code and model will be made public.
+The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+2.0 on LLaMA2-7B), instruction following (+2.9 on LLaMA2-7B) and visual instruction following (+1.4 on LLaVA1.5-7B).
+[^1] Our code and model are publicly available at https://github.com/sufenlp/MiLoRA.
 2 Preliminaries
@@
 The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order.
-Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
+Without loss of generality, we suppose $m\leq n$ to simplify the notation.
+The SVD of ${\bm{W}}$ can be reformulated as
 ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}=\sum_{i=1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$ (1)
@@
 where the singular values are in descending order ($\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}\geq 0$), and the hyperparameter $r$ is the number of minor singular values in the ${\bm{W}}_{m}$ matrix.
-The decomposition can be also reformulated in a matrix form.
+The decomposition can also be reformulated in a matrix form.
 The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principal and minor singular values, respectively.
@@
 2) Unlike vanilla LoRA which requires tuning the scaling factor $\alpha$ and the initialization hyperparameters, ours does not require any hyperparameter except the rank $r$.
-To make our method simple and training-efficient, we do not incorporate the regularization loss which imposes strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$.
+One potential variant of MiLoRA is to impose strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$ during fine-tuning.
+However, we adopt a softer approach by guiding the optimization direction solely through initialization, resulting in a method that is simpler, more training-efficient, and more flexible.
 Our experiment results show that our scheme works quite well in balancing learning from the finetuning dataset and preserving pretrained knowledge.
@@
 Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks.
+The reported metric in this table is accuracy.
 Results are cited from Liu et al.
@@
 MiLoRA 63.5 17.8 40.7
-Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH. Results are cited from Yu et al. (2023).
+Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH.
+The evaluation metric presented in this table is the Exact Match (EM) ratio.
+Results are cited from Yu et al.
+(2023).
+For each dataset, the highest performance scores among the different PEFT methods are shown in bold.
 To evaluate MiLoRA against other PEFTs, we conduct experiments on both NLP and multimodal tasks covering 18 datasets.
@@
 Models and Datasets
-We evaluate and compare MiLoRA with baselines on three different types of LLM downstream tasks.
-1.
-Commonsense reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023).
-Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018).
-The task is formulated as a multiple-choice problem.
-We report accuracy (%) for all datasets on the best checkpoint chosen by the validation set loss.
-2.
-Math reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
-We use test sets of GSM8K and MATH for evaluation and report results on the last checkpoint. For each test set, we report the exact match ratio against the ground truth.
-3.
-Instruction following: We follow the experiment setup in Wu et al.
-(2024a) and Wu et al.
-(2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023).
-Following previous works, we use Alpaca-Eval v1.0 (Li et al., 2023a) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024).
+We compare MiLoRA with baselines on three different types of LLM downstream tasks.
+[…]
+[…]
+[…]
+Method AlpacaEval 2.0 FollowBench IFEval Avg.
+LC WR WR Level1 Level2 Level3 Level4 Level5 Prompt-strict Ins-strict Prompt-loose Ins-loose
+LoRA 5.6 3.9 46.5 39.2 35.5 22.2 14.3 28.8 39.9 31.1 42.3 28.1
+PiSSA 5.6 3.9 44.2 37.1 30.6 20.9 14.3 31.6 42.7 34.6 45.7 28.3
+MiLoRA 7.1 4.4 45.1 37.6 38.2 28.5 20.8 32.4 44.8 34.8 47.7 31.0
+Table 3: Instruction following evaluation results for LLaMA2-7B on AlpacaEval 2.0, FollowBench, and IFEval.
+AlpacaEval 2.0 is evaluated using Length-controlled Win Rate (LC WR) and Win Rate (WR) against GPT-4 Turbo (OpenAI, 2023b), FollowBench adopts Hard Satisfaction Rate (HSR), and IFEval is measured by accuracy.
+Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
 Implementation Details
@@
 As math reasoning has a large training dataset, we set the rank to 64.
-Method Win-rate
-GPT-4-0613 91.2
-GPT-3.5-turbo-0613 91.8
-LLaMA2-Chat-7B 87.3
-LLaMA2-7B + LoRA 85.9
-LLaMA2-7B + PiSSA 85.6
-LLaMA2-7B + MiLoRA 87.3
-Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Results are cited from Ivison et al. (2023).
+We use the implementation of LLM-Adapters (Hu et al., 2023) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters for commonsense reasoning, the implementation of PiSSA (Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA for math reasoning and the implementation of open-instruct [^4] https://github.com/allenai/open-instruct for instruction tuning.
 Results
@@
 MiLoRA 79.2 62.1 53.3 70.6 58.7 87.9 66.1 68.3
-Table 4: Visual instruction tuning evaluation results for LLaVA1.5-7B on seven vision-language tasks. We directly use checkpoints from Liu et al. (2024a) to reproduce the results of baseline methods.
-We report results in Table 1, Table 2 and 3 for commonsense reasoning, math reasoning and instruction following, respectively.
+Table 4: Visual instruction tuning evaluation results for LLaVA1.5-7B on seven vision-language tasks.
+Accuracy serves as the metric reported in this table.
+We directly use checkpoints from Liu et al.
+(2024a) to reproduce the results of baseline methods.
+We report results in Table 1 for commonsense reasoning, Table 2 for math reasoning, and Tables 3 for instruction following, respectively.
 For commonsense reasoning, we also include the ChatGPT baseline reported in Liu et al.
-(2024b) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022).
+(2024b) as a reference, which is obtained with GPT-3.5 Turbo (OpenAI, 2023a) API using a zero-shot Chain of Thought (Wei et al., 2022).
 For math reasoning, we add the full finetuning results from Yu et al.
 (2023) as a reference.
-For instruction following, we include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2-Chat-7B from Ivison et al.
-(2023) as a reference.
 As can be seen, MiLoRA consistently surpasses all baseline methods across different datasets and LLMs, indicating that MiLoRA is a highly effective finetuning method.
-Specifically, for commonsense reasoning, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 5.4 accuracy points on LLaMA2-7B, and by 1.1 and 6.5 points on LLaMA3-8B.
-For math reasoning, MiLoRA improves over LoRA and PiSSA by 2.0 and 3.7 averaged EM points on LLaMA2-7B, respectively.
+Specifically, for commonsense reasoning, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 5.4 accuracy points on LLaMA2-7B and 1.1 and 6.5 points on LLaMA3-8B.
+For math reasoning, MiLoRA improves over LoRA and PiSSA by 2.0 and 3.7 averaged Exact Match (EM) points on LLaMA2-7B, respectively.
 However, it still underperforms full finetuning by an average of 2.5 EM score.
-This suggests that the PEFT methods still have room for improvement to fully match the performance of full finetuning.
-For instruction following, MiLoRA achieves the best results, surpassing LoRA and PiSSA by 1.4 and 2.7 in win-rate scores, respectively.
+This suggests that the PEFT methods still have room for improvement to fully match the performance of full fine-tuning.
+For instruction following, the results on AlpacaEval 2.0 indicate that PiSSA performs comparably to LoRA, while MiLoRA outperforms both by 1.5 in LC WR and 0.5 in WR, demonstrating its superior effectiveness.
+Furthermore, experiments on FollowBench and IFEval show that MiLoRA significantly surpasses both LoRA and PiSSA across these benchmarks.
+Notably, MiLoRA excels on more challenging problems (levels 4 and 5) in FollowBench, likely due to its ability to better retain pre-trained knowledge, which is crucial for handling complex instructions.
+Overall, MiLoRA surpasses LoRA and PiSSA in instruction following, achieving an average improvement of 2.9 and 2.7 points, respectively, further underscoring its advantages.
 4.2 Experiments on Vision-Language Model
@@
 5.1 Comparison with More PEFTs
+Method GSM8K Human-eval Training Cost
+rsLoRA 45.6 16.0 $\times$ 1.00
+LoRA+ 52.1 18.2 $\times$ 1.00
+DoRA 53.1 19.8 $\times$ 1.58
+AdaLoRA 50.7 17.8 $\times$ 1.07
+LoRA-GA 53.6 19.8 $\times$ 1.00
+MiLoRA 54.7 24.4 $\times$ 1.00
+Table 5: Comparison with more baselines on math and code tasks.
+Baseline GSM8K and Human-eval results are cited from Wang et al.
+(2024c).
+The evaluation metric for GSM8K is Exact Match (EM), while for Human-eval, it is Pass@1.
 To further demonstrate the effectiveness of MiLoRA, we compare MiLoRA with additional popular baselines, such as rsLoRA Kalajdzievski (2023), LoRA+ Hayou et al.
@@
 (2023), LoRA-GA Wang et al.
-(2024b).
+(2024c).
 We follow the math and code experiment settings in Wang et al.
-(2024b).
+(2024c).
 Specifically, we finetune LLaMA2-7B with MiLoRA on MetaMathQA 100K and Code-Feedback 100K for one epoch using rank $8$ and compare MiLoRA with baseline results reported in Wang et al.
-(2024b).
+(2024c).
 Table 5 shows the results.
@@
 Therefore, the training time required of MiLoRA is consistent with that of LoRA, while DoRA and AdaLoRA require additional training time overhead.
-Method GSM8K Human-eval Training Cost
-LoRA 42.1 14.8 $\times$ 1.00
-PiSSA 44.5 16.0 $\times$ 1.00
-rsLoRA 45.6 16.0 $\times$ 1.00
-LoRA+ 52.1 18.2 $\times$ 1.00
-DoRA 53.1 19.8 $\times$ 1.58
-AdaLoRA 50.7 17.8 $\times$ 1.07
-LoRA-GA 53.6 19.8 $\times$ 1.00
-MiLoRA 54.7 24.4 $\times$ 1.00
-Table 5: Comparison with more baselines on math and code tasks. Baseline GSM8K and Human-eval results are cited from Wang et al. (2024b).
 5.2 Is Minor Singular Component Adaptation Important?
+Method Principal Random Minor
+GSM8K 60.7 63.2 64.0
+MATH 14.6 15.5 16.1
+Table 6: Performance of MiLoRA when initializing with principal, random sampled, and minor singular components.
 To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model.
@@
 This underscores the importance of adapting the minor singular components during the finetuning process.
-Concurrent with our work, Meng et al.
-(2024) propose PiSSA, a similar approach to MiLoRA but adapts the principal singular components.
-We will provide a detailed comparison between MiLoRA and PiSSA in Section 5.5.
-Method principal Random Minor
-GSM8K 60.7 63.2 64.0
-MATH 14.6 15.5 16.1
-Table 6: Performance of MiLoRA when initializing with principal, random sampled, and minor singular components.
+$\left\|{\bm{W}}\right\|_{F}=97.36$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
+$\left\|{\bm{U}}^{\top}{\bm{W}}{\bm{V}}\right\|_{F}=$ 43.64 1.50 1.82 17.57 1.86
+$\left\|{\bm{U}}^{\top}\Delta{\bm{W}}{\bm{V}}\right\|_{F}=$ 68.18 55.79 44.95
+Amplification factor - - 37.46 3.18 24.17
+Table 7: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs.
+The weight matrices are taken from the query projection of the 16th layer of LLaMA2-7B.
+Figure 2: Left: The subspace similarity averaged by all modules in a layer. Right: The subspace similarity averaged by all layers for a specific module type.
 5.3 How does the Matrix Update $\Delta W$ Compare to $W$?
@@
 As a comparison, we also compute ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ by replacing ${\bm{U}}$, ${\bm{V}}$ with the top $r$ singular vectors of ${\bm{W}}$ or a random matrix.
-We use the Frobenius norm as the measurement of magnitude.
 The results are shown in Table 7.
 Following Hu et al.
-(2021), we use the query weight in the middle layer (16th) of the model for analysis.
+(2021), we use the query weight in the middle layer (16th) of the model for analysis and use the Frobenius norm as the measurement of magnitude.
 We make two key observations.
@@
 This indicates that MiLoRA has a reduced impact on the features which already present in ${\bm{W}}$.
-See Table 13 of Appendix B for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
-$\left\|{\bm{W}}\right\|_{F}=97.36$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA
-$\left\|{\bm{U}}^{\top}{\bm{W}}{\bm{V}}\right\|_{F}=$ 43.64 1.50 1.82 17.57 1.86
-$\left\|{\bm{U}}^{\top}\Delta{\bm{W}}{\bm{V}}\right\|_{F}=$ 68.18 55.79 44.95
-Amplification factor - - 37.46 3.18 24.17
-Table 7: The Frobenius norm of ${\bm{U}}^{\top}{\bm{W}}{\bm{V}}$ where ${\bm{U}}$ and ${\bm{V}}$ are the left/right top $r$ singular vectors of either ${\bm{W}}$, a random matrix, or $\Delta{\bm{W}}$ of PEFTs.
-The weight matrices are taken from the query projection of the 16th layer of LLaMA2-7B.
-Figure 2: Left: The subspace similarity averaged by all modules in a layer. Right: The subspace similarity averaged by all layers for a specific module type.
+See Table 14 of Appendix B.2 for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
 To further analyze the characteristics of $\Delta{\bm{W}}$ for different methods, we measure the normalized subspace similarity (Hu et al., 2021) between the subspaces spanned by top $r$ singular vectors of $\Delta{\bm{W}}$ and the subspace spanned by top $r$ singular vectors, bottom $r$ singular vectors, and a random $r$ singular vectors of ${\bm{W}}$.
 We define the normalized similarity metric based on the Grassmann distance following Hu et al.
-(2021) (See Appendix A.1 for more details).
-The results are presented in Figure 2.
-From the left figure, it is evident that PiSSA focuses on optimizing features associated with the top singular vectors, whereas LoRA and MiLoRA do not significantly optimize these features.
-MiLoRA primarily optimizes features related to the bottom singular vectors.
-The right figures further show that MiLoRA places greater emphasis on features associated with the bottom singular vectors of the MLP layers.
+(2021) (See Appendix A.1 for more details) and report the results in Figure 2.
+From the left subfigure, it is clear that PiSSA learns in directions closely aligned with the top singular vectors of the pretrained weights, while LoRA and MiLoRA do not significantly optimize in these directions.
+Specifically, MiLoRA focuses on optimizing directions associated with the bottom singular vectors of the pretrained weights.
+The right subfigure further shows that weight matrices in both MLP and self attention modules of MiLoRA learn in directions more closely aligned with the bottom singular vectors, especially the MLP modules.
 5.4 MiLoRA Forgets Less than Baselines
@@
 In contrast, our method MiLoRA aims to adapt to new tasks while maximally retaining the base model’s knowledge.
-To achieve this, we instead finetune the minor singular components of the weight matrices, which are less important for the weight matrices.
+To achieve this, we instead finetune the minor singular components of the weight matrices, which are less important for the pretrained weight matrices.
 Performance: The PiSSA paper claims that PiSSA outperforms both LoRA and adapting minor singular components.
@@
 However, they introduce additional computational costs during inference.
-LoRA-based methods (Hu et al., 2021; Liu et al., 2024b; Zhang et al., 2023) model the changes of selected linear weights as low-rank matrices.
+LoRA-based methods (Hu et al., 2021; Liu et al., 2024b; Zhang et al., 2023; Wang et al., 2024a) model the changes of selected linear weights as low-rank matrices.
 During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen.
@@
 Concurrent with our work, PISSA (Meng et al., 2024) and LoRA-GA Wang et al.
-(2024b) propose to better approximate full finetuning by only updating the principal singular components or aligning the gradients of low-rank updates to that of full finetuning.
+(2024c) propose to better approximate full finetuning by only updating the principal singular components or aligning the gradients of low-rank updates to that of full finetuning.
 7 Conclusion
@@
 Limitations
-Due to limited computational resources, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks following previous work Liu et al.
-(2024b); Wu et al.
+Due to limited computational resources, we primarily evaluate the effectiveness of MiLoRA on LLaMA-family LLMs and LLaVA-1.5, focusing on tasks such as commonsense reasoning, math reasoning, and instruction following, in line with prior work Liu et al.
+(2024b); Meng et al.
+(2024); Wu et al.
 (2024b).
@@
 (2023); Liu et al.
-(2024b) instead of exhaustive hyperparameter search for each task.
+(2024a) instead of an exhaustive hyperparameter search for each task.
 We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
+Acknowledgements
+This project was supported by National Natural Science Foundation of China (No.
+62306132, No.
+62106138) and Guangdong Basic and Applied Basic Research Foundation.
+We thank the anonymous reviewers for their insightful feedback on this work.
 Appendix A Detailed Experiment Setups
@@
 Table 12: The detailed hyperparameter configurations used by PiSSAMeng et al. (2024).
-A.4 More Implementation Details
-We use the implementation of LLM-Adapters (Hu et al., 2023) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters for commonsense reasoning, the implementation of PiSSA (Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA for math reasoning and the implementation of open-instruct [^4] https://github.com/allenai/open-instruct for instruction tuning.
+A.4 Hyperparameter Setup for Vision-Language Model
 For hyperparameters of finetuning the vision-language model, we follow the LoRA configuration provided by Liu et al.
@@
 Appendix B More Experiment Results
+B.1 Experiments on Qwen2.5-7B
+Method GSM8K MATH Avg.
+LoRA 81.8 49.3 65.5
+PiSSA 81.7 45.4 63.6
+MiLoRA 85.5 48.7 67.1
+Table 13: Math reasoning evaluation results for Qwen2.5-7B on GSM8K and MATH.
+To enhance the credibility of our experiments, we evaluate MiLoRA using Qwen2.5-7B on the MetaMathQA-100K dataset.
+As shown in Table 13, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 3.5 EM scores, respectively, demonstrating its effectiveness on the Qwen backbone.
+Additionally, we conduct experiments on LLMs such as LLaMA2 and LLaMA3 (Section 4.1), as well as vision-language models like LLaVA-1.5 (Section 4.2).
+Collectively, these results confirm the effectiveness of MiLoRA across diverse model architectures.
+B.2 Complementary Weighet Metrics Analysis
 $\left\|{\bm{W}}\right\|_{F}=123.0$ ${\bm{W}}$ Random LoRA PiSSA MiLoRA